OpenStudy (anonymous):

a student can type 180words in 3mins.and 368 words 7mins.How many words can the student type in 10min?? A.470 B.548 C.509 D.587

OpenStudy (lgbasallote):

try finding how many words a student can type in 1 minute

OpenStudy (lgbasallote):

i thought so too @ParthKohli but look at 7 minutes

OpenStudy (lgbasallote):

how about adding 7 mins and 3 mins?

OpenStudy (amistre64):

umm, 7+3 = 10

Parth (parthkohli):

Oh yeah it'd be multiply that by 7.

OpenStudy (lgbasallote):

like i said lol :p

OpenStudy (amistre64):

im getting too slow at the keyboard in my old age

Parth (parthkohli):

Sorry :(

Parth (parthkohli):

Weird how all 99's are answering the question.

OpenStudy (anonymous):

ok so is someone going to help me

OpenStudy (amistre64):

its not that hard to get to a 99 ...

OpenStudy (lgbasallote):

we just did?

OpenStudy (amistre64):

does sherri define help as: just give me the answer?

Parth (parthkohli):

7 + 3 = 10 This is how it's done

OpenStudy (lgbasallote):

a student can type 180 words in 3 minutes 368 words in 7 minutes 7 minutes + 3 minutes = 10 minutes therefore add how many words the student types in 3 minutes + how many words the student types in 7 minutes

Parth (parthkohli):

180 + 368 = ?

OpenStudy (anonymous):

Seems to be a non-linear function of time.

OpenStudy (anonymous):

Exactly. The ratio isnt constant. I dont think you can just add it up.

OpenStudy (amistre64):

this doesnt seem to be a "fit a curve" question. it gives a 3 minute and a 7 minute value and asks for a 10 minute value

Parth (parthkohli):

We have to find the rate of change here Iguess

OpenStudy (anonymous):

I know, Amistre. I was being tongue-in-cheek.

OpenStudy (amistre64):

:) i had considered using 0,0 and determining a sqrt or ln to fit tho

OpenStudy (amistre64):

or parabola depending on where the points fall

OpenStudy (amistre64):

a bajillion is my guess tho

OpenStudy (anonymous):

Hmm.. Maybe we can solve it generally?

OpenStudy (anonymous):

470?

OpenStudy (anonymous):

Actually, it's 42. It's definitely 42.

OpenStudy (lgbasallote):

cant we just agree on the simple imaginary solution :p lol

OpenStudy (anonymous):

Hmm.. Maybe you're right, lgba.